OpenStudy (anonymous):

WILL GIVE MEDAL! :) Simplify (2√5+3√7)^2 show your work. justify each step.

OpenStudy (studygurl14):

\(\Large (2\sqrt{5}+3\sqrt{7})^2=(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})=?\)

OpenStudy (studygurl14):

Here's a hint: \(\Large (a+b)(a+b)=a^2+2ab+b^2\)

OpenStudy (anonymous):

Wait how do i multiply them together? Do i use the Pythagoreum Theorm?

OpenStudy (studygurl14):

No, not the Pythagorean theorem. Look at the hint I gave you...In this situation, \(\large a=2\sqrt{5}\) and \(\large b=3\sqrt{7}\)

OpenStudy (anonymous):

FOIL: first, outside, inside, last

OpenStudy (studygurl14):

^

OpenStudy (anonymous):

Will you help me square it? I don't think i know how to properly square radical expressions.

OpenStudy (studygurl14):

If you use the FOIL method...